each of X contains Y boxes. each box contain Z marble .how many marbles are ther in all
 

Grade:12th pass

2 Answers

Kriti
2 years ago
Each X has Y boxes 
So, Total number of boxes =X*Y
Each box has Z marbles 
So Total number of marbles= X*Y*Z
Debarati
2 years ago
Number of boxes = x*y = xy
Total number of marbles = x*y*z = xyz
Therefore,the total number of marbles altogether is xyz.